Find how fast (a) the volume, (b) the surface area, (c) the diagonal, of
a cube increases when the length of the edge increases
1
Expert's answer
2020-12-14T07:29:45-0500

A) Let "V_0" – initial the volume and "V_1" - the volume when the length of the edge increases, "a_0" – initial length of the edge, "a_1" – new length of the edge, then "V_1=\frac{a_1^3}{a_0^3}\times V_0" .

B) Let "S_0" - initial the surface area and "S_1" - the surface area when the length of the edge increases, "a_0" – initial length of the edge, "a_1" – new length of the edge, then "S_1=\frac{a_1^2}{a_0^2}\times S_0" .

C) Similar to the previous example "d_1=\frac{a_1^2}{a_0^2}\times d_0" .